# Demystifying amortized causal discovery with transformers

### v1

@@ Abstract
Supervised learning approaches for causal discovery from observational data often achieve competitive performance despite seemingly avoiding explicit assumptions that traditional methods make for identifiability. In this work, we investigate CSIvA [1], a transformer-based model promising to train on synthetic data and transfer to real data. First, we bridge the gap with existing identifiability theory and show that constraints on the training data distribution implicitly define a prior on the test observations. Consistent with classical approaches, good performance is achieved when we have a good prior on the test data, and the underlying model is identifiable. At the same time, we find new trade-offs. Training on datasets generated from different classes of causal models, unambiguously identifiable in isolation, improves the test generalization. Performance is still guaranteed, as the ambiguous cases resulting from the mixture of identifiable causal models are unlikely to occur (which we formally prove). Overall, our study finds that amortized causal discovery still needs to obey identifiability theory, but it also differs from classical methods in how the assumptions are formulated, trading more reliance on assumptions on the noise type for fewer hypotheses on the mechanisms.
@@ 1 Introduction
Causal discovery aims to uncover the underlying causal relationships between variables of a system from pure observations, which is crucial for answering interventional and counterfactual queries when experimentation is impractical or unfeasible [2, 3, 4]. Unfortunately, causal discovery is inherently ill-posed [5]: unique identification of causal directions requires restrictive assumptions on the class of structural causal models (SCMs) that generated the data [6, 7, 8]. These theoretical limitations often render existing methods inapplicable, as the underlying assumptions are usually untestable or difficult to verify in practice [9].
@@
Recently, supervised learning algorithms trained on synthetic data have been proposed to overcome the need for specific hypotheses, which restrains the application of classical causal discovery methods to real-world problems [1, 10, 11, 12, 13]. Seminal work from Lopez-Paz et al. [10] argues that this learning-based approach to causal discovery would allow dealing with complex data-generating processes and would greatly reduce the need for explicitly crafting identifiability conditions a-priori: despite this ambitious goal, the output of these methods is generally considered unreliable, as no theoretical guarantee is provided. A pair of non-identifiable structural causal models can be associated with different causal graphs $\mathcal{G}\neq\tilde{\mathcal{G}}$, while entailing the same joint distribution $p$ on the system’s variables. It is thus unclear how a learning algorithm presented with observational data generated from $p$ would be able to overcome these theoretical limits and correctly identify a unique causal structure. However, the available empirical evidence seems not to care about impossibility results, as these methods yield surprising generalization results on several synthetic benchmarks. Our work aims to bridge this gap by studying the performance of a transformer architecture for causal discovery through the lens of the theory of identifiability from observational data. Specifically, we analyze the CSIvA (Causal Structure Induction via Attention) model for causal discovery [1], focusing on bivariate graphs, as they offer a controlled yet non-trivial setting for the investigation. As our starting point, we provide closed-form examples that identify the limitations of CSIvA in recovering causal structures of linear non-Gaussian and nonlinear additive noise models, which are notably identifiable, and demonstrate the expected failures through empirical evidence. These findings suggest that the class of structural causal models that can be identified by CSIvA is inherently dependent on the specific class of SCMs observed during training. Thus, the need for restrictive hypotheses on the data-generating process is intrinsic to causal discovery, both in the traditional and modern learning-based approaches: assumptions on the test distribution either are posited when selecting the algorithm (traditional methods) or in the choice of the training data (learning-based methods). To address this limitation, we theoretically and empirically analyze when training CSIvA on datasets generated by multiple identifiable SCMs with different structural assumptions improves its generalization at test time. In summary:
@@
  • We show that the class of structural causal models that CSIvA can identify is defined by the class of SCMs observed through samples during the training. We reinforce the notion that identifiability in causal discovery inherently requires assumptions, which must be encoded in the training data in the case of learning-based approaches.
@@
  • To overcome this limitation, we study the benefits of CSIvA training on mixtures of causal models. We analyze when algorithms learned on multiple models are expected to identify broad classes of SCMs (unlike many classical methods). Empirically, we show that training on samples generated by multiple identifiable causal models with different assumptions on mechanisms and noise distribution results in significantly improved generalization abilities.
Closely related works and their relation with CSIvA.
@@
In this paper, we study amortized inference of causal graphs, i.e. optimization of an inference model to directly predict a causal structure from newly provided data. This is the first work that attempts to understand the connection between identifiability theory and amortized inference, while several algorithms have been proposed. In the context of purely observational data, Lopez-Paz et al. [10] defines a distribution regression problem [14] mapping the kernel mean embedding of the data distribution to a causal graph, while Li et al. [11] relies on equivariant neural network architectures. More recently, Lippe et al. [12] and Lorch et al. [13] proposed learning on interventional data, in addition to observations (in the same spirit as CSIvA). Despite different algorithmic implementations, the target object of estimation of most of these methods is the distribution over the space of all possible graphs, conditional on the input dataset (similarly, the ENCO algorithm in Lippe et al. [12] models the conditional distribution of individual edges). This justifies our choice of restricting our study to the CSIvA architecture (despite this being a clear limitation), as in the infinite observational sample limit, these methods approximate the same distribution. Methods necessarily requiring interventional data [15, 16, 17], and learning-based algorithms unsuitable for amortized inference [18, 19, 20, 21, 22] are out of the scope of this work.
@@ 2.1 Causal discovery from observational data
with $f_{2,i}$ invertible map and mutually independent noise terms. When $f_{2,i}$ is a nonlinear function, the post-ANM amounts to the identifiable post-nonlinear model (PNL) [8]. When $f_{2,i}$ is the identity function and $f_{1,i}$ nonlinear, it simplifies to the nonlinear additive noise model (ANM)[7, 23], which is known to be identifiable, and is described by the set of structural equations:
@@
If, additionally, we restrict the mechanisms $f_{1,i}$ to be linear and the noise terms $N_{i}$ to a non-Gaussian distribution, we recover the identifiable linear non-Gaussian additive model or LiNGAM [6]:
@@ 2.2 Motivation and problem definition
Causal discovery from observational data relies on specific assumptions, which can be challenging to verify in practice [9]. To address this, recent methods leverage supervised learning for the amortized inference of causal graphs [1, 10, 11, 12, 13, 16, 24], optimizing an inference model to directly predict a causal structure from a provided dataset. While these approaches aim to reduce reliance on explicit identifiability assumptions, they often lack a clear connection to the existing causal discovery theory, making their outputs generally unreliable. We illustrate this limitation through an example.
@@
We consider the CSIvA transformer architecture proposed by Ke et al. [1], which can learn a map from observational data to a causal graph. The authors of the paper show that, in the infinite sample regime, the CSIvA architecture exactly approximates the conditional distribution $p(\cdot|\mathcal{D})$ over the space of possible graphs, given a dataset $\mathcal{D}$. Identifiability theory in causal discovery tells us that if the class of structural causal models that generated the observations is sufficiently constrained, then there is only one graph that can fit the data within that class. For example, consider the case of a dataset that is known to be generated by a nonlinear additive noise model, and let $p(\cdot|\mathcal{D},\textnormal{ANM})$ be the conditional distribution that incorporates this prior knowledge on the SCM: then $p(\cdot|\mathcal{D},\textnormal{ANM})$ concentrates all the mass on a single point $\mathcal{G}^{*}$, the true graph underlying the $\mathcal{D}$ observations. Instead, in the absence of restrictions on the structural causal model, all the graphs in a Markov equivalence class are equally likely to be the correct solution given the data. Hence, $p(\cdot|\mathcal{D})$, the distribution learned by CSIvA, assigns equal probability to each graph in the Markov equivalence class of $\mathcal{G}^{*}$.
@@
Our arguments of Example 1 are valid for all learning methods that approximate the conditional distribution over the space of graphs given the input data [1, 10, 11, 12, 13], and suggest that these algorithms are at most informative about the equivalence class of the causal graph underlying the observations. However, the available empirical evidence does not seem to highlight these limitations, as in practice these methods can infer the true causal DAG on several synthetic benchmarks. Thus, further investigation is necessary if we want to rely on their output in any meaningful sense. In this work, we analyze these "black-box" approaches through the lens of established theory of causal discovery from observational data (causal inference often lacks experimental data, which we do not consider). We study in detail the CSIvA architecture [1] (see Appendix A), a variation of the transformer neural network [25] for the supervised learning of algorithms for amortized causal discovery. This model is optimized via maximum likelihood estimation, i.e. finding $\Theta$ that minimizes $-\mathbf{E}_{\mathcal{G},\mathcal{D}}[\ln\hat{p}(\mathcal{G}|\mathcal{D};\Theta)]$, where $\hat{p}(\mathcal{G}|\mathcal{D};\Theta)$ is the conditional distribution of a graph $\mathcal{G}$ given a dataset $\mathcal{D}$ parametrized by $\Theta$. We limit the analysis to CSIvA as it is a simple yet competitive end-to-end approach to learning causal models. While this is clearly a limitation of the paper, our theoretical and empirical conclusions exemplify both the role of theoretical identifiability in modern approaches and the new opportunities they provide. Additionally, it fits well within a line of works arguing that specifically transformers can learn causal concepts [26, 27, 28] and identify different assumptions in context [29].
@@ 3 Experimental results through the lens of theory
In this section, we present a comprehensive analysis of causal discovery with transformers and its relation to the theoretical boundaries of causal discovery from observational data. We show that suitable assumptions must be encoded in the training distribution to ensure the identifiability of the test data, and we additionally study the effectiveness of training on mixtures of causal models to overcome these limitations, improving generalization abilities.
@@ 3.1 Experimental design
We concentrate our research on causal models of two variables, causally related according to one of the two graphs $X\rightarrow Y$, $Y\rightarrow X$. Bivariate models are the simplest non-trivial setting with a well-known theory of causality inference [7, 8, 23], but also amenable to manipulation. This allows for comprehensive training and analysis of diverse SCMs and facilitates a clear interpretation of the results.
@@
Unless otherwise specified, in our experiments we train CSIvA on a sample of $15000$ synthetically generated datasets, consisting of $1500$ i.i.d. observations. Each dataset is generated according to a single class of SCMs, defined by the mechanism type and the noise terms distribution. The coefficients of the linear mechanisms are sampled in the range $[-3,-0.5]\cup[0.5,3]$, removing small coefficients to avoid close-to-unfaithful effects [30]. Nonlinear mechanisms are parametrized according to a neural network with random weights, a strategy commonly adopted in the literature of causal discovery [1, 9]. The post-nonlinearity of the PNL model consists of a simple map $z\mapsto z^{3}$. Noise terms are sampled from common distributions and a randomly generated density that we call mlp, previously adopted in Montagna et al. [9], defined by a standard Gaussian transformed by a multilayer perceptron (MLP) (Appendix B.2). We name these datasets mechanism-noise to refer to their underlying causal model. For example, data sampled from a nonlinear ANM with Gaussian noise are named nonlinear-gaussian. More details on the synthetic data generation schema are found in Appendix B.2. All data are standardized by their empirical variance to remove opportunities to learn shortcuts [31, 32, 33].
@@
As our metric we use the structural Hamming distance (SHD), which is the number of edge removals, insertions or flips to transform one graph to another. In the context of bivariate causal graphs with a single edge, this is simply an error count, so correct inference corresponds to $\textnormal{SHD}=0$, and an incorrect prediction gives $\textnormal{SHD}=1$. Additionally, we define a reference random baseline, which assigns a causal direction according to a fair coin, achieving $\textnormal{SHD}=0.5$ in expectation. Each architecture we analyze in the experiments is trained $3$ times, with different parameter initialization and training samples: the SHD presented in the plots is the average of each of the $3$ models on $1500$ distinct test datasets of $1500$ points each, and the error bars are $95\%$ confidence intervals.

### v2

@@ Abstract
Supervised learning for causal discovery from observational data often achieves competitive performance despite seemingly avoiding the explicit assumptions that traditional methods require for identifiability. In this work, we analyze CSIvA (Ke et al., 2023b) on bivariate causal models, a transformer architecture for amortized inference promising to train on synthetic data and transfer to real ones. First, we bridge the gap with identifiability theory, showing that the training distribution implicitly defines a prior on the causal model of the test observations: consistent with classical approaches, good performance is achieved when we have a good prior on the test data, and the underlying model is identifiable. Second, we find that CSIvA can not generalize to classes of causal models unseen during training: to overcome this limitation, we show that learning on datasets generated from different types of causal models, unambiguously identifiable in isolation, improves the test generalization. We analyze this empirical evidence with theory, illustrating that the ambiguous cases resulting from the mixture of identifiable causal models are unlikely to occur. Overall, we find that amortized causal discovery still adheres to identifiability theory, violating the previous hypothesis from Lopez-Paz et al. (2015) that supervised learning methods could overcome its restrictions.
@@ 1 Introduction
Causal discovery aims to uncover the underlying causal relationships between variables of a system from pure observations, which is crucial for answering interventional and counterfactual queries when experimentation is impractical or unfeasible (Peters et al., 2017, Pearl, 2009, Spirtes, 2010). Unfortunately, causal discovery is inherently ill-posed (Glymour et al., 2019): unique identification of causal directions requires restrictive assumptions on the class of structural causal models (SCMs) that generated the data (Shimizu et al., 2006, Hoyer et al., 2008, Zhang & Hyvärinen, 2009). These theoretical limitations often render existing methods inapplicable, as the underlying assumptions are usually untestable or difficult to verify in practice (Montagna et al., 2023a).
@@
Recently, supervised learning algorithms trained on synthetic data have been proposed to overcome the need for specific hypotheses, which restrains the application of classical causal discovery methods to real-world problems (Ke et al., 2023b, Lopez-Paz et al., 2015, Li et al., 2020, Lippe et al., 2022, Lorch et al., 2022). Seminal work from Lopez-Paz et al. (2015) argues that this learning-based approach to causal discovery would allow dealing with complex data-generating processes and would greatly reduce the need for explicitly crafting identifiability conditions a-priori: despite this ambitious goal, the output of these methods is generally considered unreliable, as no theoretical guarantee is provided. A pair of non-identifiable structural causal models can be associated with different causal directed acyclic graphs (DAGs) $\mathcal{G}\neq\tilde{\mathcal{G}}$, while entailing the same joint distribution $p$ on the system’s variables. It is thus unclear how a learning algorithm presented with observational data generated from $p$ would be able to overcome these theoretical limits and correctly identify a unique causal structure. However, the available empirical evidence seems not to reflect impossibility results, as these methods yield surprising generalization abilities on several synthetic benchmarks. Our work aims to bridge this gap by studying the performance of a transformer architecture for causal discovery through the lens of the theory of identifiability from observational data. Specifically, we analyze the CSIvA (Causal Structure Induction via Attention) model for causal discovery (Ke et al., 2023b), focusing on bivariate graphs, as they offer a controlled yet non-trivial setting for the investigation. As our starting point, we provide closed-form examples that identify the limitations of CSIvA in recovering causal structures of linear non-Gaussian and nonlinear additive noise models, which are notably identifiable, and demonstrate the expected failures through empirical evidence. These findings suggest that the class of structural causal models that can be identified by CSIvA is inherently dependent on the specific class of SCMs observed during training. Thus, the need for restrictive hypotheses on the data-generating process is intrinsic to causal discovery, both in the traditional and modern learning-based approaches: assumptions on the test distribution either are posited when selecting the algorithm (traditional methods) or in the choice of the training data (learning-based methods). To address this limitation, we theoretically and empirically analyze when training CSIvA on datasets generated by multiple identifiable SCMs with different structural assumptions improves its generalization at test time. Our experimental findings are based on the analysis of $\sim 1$M test runs. In summary:
@@
  • We show that the class of structural causal models that CSIvA can identify is defined by the class of SCMs observed through samples during the training. We reinforce the notion that identifiability in causal discovery inherently requires assumptions, which must be encoded in the training data in the case of learning algorithms for amortized inference: this violates a previous hypothesis in Lopez-Paz et al. (2015), which suggests that these methods could exceed the boundaries of identifiability.
  • We empirically show that CSIvA is expected to fail to generalize on datasets generated by structural causal models characterized by mechanism types or noise terms distributions unseen during training. While this appears as a significant limit of amortized causal discovery, systematic analysis has been disregarded by previous work in the literature.
@@
  • To mitigate this limitation, we study the benefits of CSIvA training on mixtures of causal models. We analyze when algorithms learned on multiple models are expected to identify broad classes of SCMs (unlike many classical methods). Empirically, we show that training on samples generated by multiple identifiable causal models with different assumptions on mechanisms and noise distribution results in significantly improved generalization abilities.
2 Related works
In this paper, we study amortized inference of bivariate causal graphs, i.e. supervised optimization of an inference model to directly predict a causal structure from newly provided data. In particular, this is the first work that draws a connection between identifiability theory and amortized inference of causal DAGs. Dai et al. (2023) studies supervised learning of the graph skeleton, limiting its analysis to the role of identifiability of unshielded triplets. Several algorithms have instead been proposed.
Algorithms for amortized inference closely related to CSIvA.
@@ 2 Related works
In the context of purely observational data, Lopez-Paz et al. (2015) defines a classification problem mapping the kernel mean embedding of the data distribution to a causal graph, while Li et al. (2020) relies on equivariant neural network architectures. More recently, Lippe et al. (2022) and Lorch et al. (2022) proposed learning on interventional data, in addition to observations (in the same spirit as CSIvA). Despite different algorithmic implementations, the target object of estimation of most of these methods is the distribution over the space of all possible graphs, conditional on the input dataset (similarly, the ENCO algorithm in Lippe et al. (2022) models the conditional distribution of individual edges). This justifies our choice of restricting our study to the CSIvA architecture (despite this being a clear limitation), as in the infinite observational sample limit, these methods approximate the same distribution.
Other learning-based algorithms for causal discovery.
Out of the scope of this work, there are methods that necessarily require interventional data (Brouillard et al., 2020, Ke et al., 2023a, Scherrer et al., 2022), and learning-based algorithms unsuitable for amortized inference (Lachapelle et al., 2020, Ng et al., 2020, Zheng et al., 2018, Zhang et al., 2022, Bello et al., 2022).
Differences with Lopez-Paz et al. (2015).
Before moving forward, we remark on the main differences between our paper and Lopez-Paz et al. (2015), as both works concentrate on supervised learning for the inference of bivariate causal graphs. Lopez-Paz et al. (2015) frames causal discovery as a classification problem, where the goal is estimating and mapping the kernel mean embeddings of the distribution of the observed data to the correct causal order (assuming a causal relation is in place). Building on the theory of reproducing kernel Hilbert spaces, they provide finite sample learning rates. In particular, their study assumes observations generated by identifiable causal models. In contrast, we aim to (i) empirically investigate what conditions enable identifiability in amortized causal discovery (ii) theoretically and empirically investigate how to exploit the known identifiability results to train algorithms with improved test generalization.
@@ 3.1 Causal discovery from observational data
with $f_{2,i}$ invertible map and mutually independent noise terms. When $f_{2,i}$ is a nonlinear function, the post-ANM amounts to the identifiable post-nonlinear model (PNL) (Zhang & Hyvärinen, 2009). When $f_{2,i}$ is the identity function and $f_{1,i}$ nonlinear, it simplifies to the nonlinear additive noise model (ANM)(Hoyer et al., 2008, Peters et al., 2014), which is known to be identifiable, and is described by the set of structural equations:
@@
If, additionally, we restrict the mechanisms $f_{1,i}$ to be linear and the noise terms $N_{i}$ to a non-Gaussian distribution, we recover the identifiable linear non-Gaussian additive model or LiNGAM (Shimizu et al., 2006):
@@ 3.2 Motivation and problem definition
Causal discovery from observational data relies on specific assumptions, which can be challenging to verify in practice (Montagna et al., 2023a). To address this, recent methods leverage supervised learning for the amortized inference of causal graphs (or simply amortized causal discovery), i.e. optimization of an inference model to directly predict a causal structure from newly provided data (Lopez-Paz et al., 2015, Li et al., 2020, Lippe et al., 2022, Lorch et al., 2022, Ke et al., 2023a, Löwe et al., 2020). While these approaches also aim to reduce reliance on explicit identifiability assumptions, they often lack a clear connection to the existing causal discovery theory, making their outputs generally unreliable. We illustrate this limitation through an example.
@@
We consider the CSIvA transformer architecture proposed by Ke et al. (2023b), which can learn a map from observational data to a causal graph. The authors of the paper show that, in the infinite sample regime, the CSIvA architecture exactly approximates the conditional distribution $p(\cdot|\mathcal{D})$ over the space of possible graphs, given a dataset $\mathcal{D}$. Identifiability theory in causal discovery tells us that if the class of structural causal models that generated the observations is sufficiently constrained, then there is only one graph that can fit the data within that class. For example, consider the case of a dataset that is known to be generated by a nonlinear additive noise model, and let $p(\cdot|\mathcal{D},\textnormal{ANM})$ be the conditional distribution that incorporates this prior knowledge on the SCM: then $p(\cdot|\mathcal{D},\textnormal{ANM})$ concentrates all the mass on a single point $\mathcal{G}^{*}$, the true graph underlying the $\mathcal{D}$ observations. Instead, in the absence of restrictions on the structural causal model, all the graphs in a Markov equivalence class are equally likely to be the correct solution given the data. Hence, $p(\cdot|\mathcal{D})$, the distribution learned by CSIvA, assigns equal probability to each graph in the Markov equivalence class of $\mathcal{G}^{*}$.
@@
Our arguments of Example 1 are valid for all learning methods that approximate the conditional distribution over the space of graphs given the input data (Ke et al., 2023b, Lopez-Paz et al., 2015, Li et al., 2020, Lippe et al., 2022, Lorch et al., 2022), and suggest that these algorithms are at most informative about the equivalence class of the causal graph underlying the observations. However, the available empirical evidence does not seem to highlight these limitations, as in practice these methods can infer the true causal DAG on several synthetic benchmarks. Thus, further investigation is necessary if we want to rely on their output in any meaningful sense. In this work, we analyze these "black-box" approaches through the lens of established theory of causal discovery from observational data (causal inference often lacks experimental data, which we do not consider). We study in detail the CSIvA architecture (Ke et al., 2023b) (see Appendix A), a variation of the transformer neural network (Vaswani et al., 2017) for the supervised learning of algorithms for amortized causal discovery. This model is optimized via maximum likelihood estimation, i.e. finding $\Theta$ that minimizes $-\mathbf{E}_{\mathcal{G},\mathcal{D}}[\ln\hat{p}(\mathcal{G}|\mathcal{D};\Theta)]$, where $\hat{p}(\mathcal{G}|\mathcal{D};\Theta)$ is the conditional distribution of a graph $\mathcal{G}$ given a dataset $\mathcal{D}$ parametrized by $\Theta$. We limit the analysis to CSIvA as it is a simple yet competitive end-to-end approach to learning causal models. While this is clearly a limitation of the paper, our theoretical and empirical conclusions exemplify both the role of theoretical identifiability in modern approaches and the new opportunities they provide. Additionally, it fits well within a line of works arguing that specifically transformers can learn causal concepts Jin et al. (2024), Zhang et al. (2024), Scetbon et al. (2024) and can be explicitly trained to identify different assumptions in context (Gupta et al., 2023).
@@ 4 Experimental results through the lens of theory
In this section, we present a comprehensive analysis of bivariate causal discovery with transformers and its relation to the theoretical boundaries of causal discovery from observational data. We show that suitable assumptions must be encoded in the training distribution to ensure the identifiability of the test data, and we additionally study the effectiveness of training on mixtures of causal models to overcome these limitations, improving generalization abilities. In Appendix C we discuss how our findings can be extended to the case of multivariate causal models.
@@ 4.1 Experimental design
We concentrate our research on causal models of two variables, causally related according to one of the two graphs $X\rightarrow Y$, $Y\rightarrow X$. Bivariate models are the simplest non-trivial setting with a well-known theory of causality inference (Hoyer et al., 2008, Zhang & Hyvärinen, 2009, Peters et al., 2014), but also amenable to manipulation. This allows for comprehensive training and analysis of diverse SCMs and facilitates a clear interpretation of the results.
@@
Unless otherwise specified, in our experiments we train CSIvA on a sample of $15000$ synthetically generated datasets, consisting of $1500$ i.i.d. observations. Each dataset is generated according to a single class of SCMs, defined by the mechanism type and the noise terms distribution. The coefficients of the linear mechanisms are sampled in the range $[-3,-0.5]\cup[0.5,3]$, removing small coefficients to avoid close-to-unfaithful effects (Uhler et al., 2012). Nonlinear mechanisms are parametrized according to a neural network with random weights, a strategy commonly adopted in the literature of causal discovery (see Appendix B.2; alternatively, we provide experiments on data generated simulating nonlinear mechanisms by sampling from a Gaussian process, as described in Appendix D.7). The post-nonlinearity of the PNL model consists of a simple map $z\mapsto z^{3}$. Noise terms are sampled from common distributions and a randomly generated density that we call mlp, previously adopted in Montagna et al. (2023a), defined by a standard Gaussian transformed by a multilayer perceptron (MLP) (Appendix B.2). We name these datasets mechanism-noise to refer to their underlying causal model. For example, data sampled from a nonlinear ANM with Gaussian noise are named nonlinear-gaussian. More details on the synthetic data generation schema are found in Appendix B.2. All data are standardized by their empirical variance to remove opportunities to learn shortcuts (Geirhos et al., 2020, Reisach et al., 2021, Montagna et al., 2023b).
@@
As our metric we use the structural Hamming distance (SHD), which is the number of edge removals, insertions or flips required to transform the predicted graph to the ground-truth. In the context of bivariate causal graphs with a single edge, this is simply an error count, so correct inference corresponds to $\textnormal{SHD}=0$, and an incorrect prediction gives $\textnormal{SHD}=1$. Additionally, we define a reference random baseline, which assigns a causal direction according to a fair coin, achieving $\textnormal{SHD}=0.5$ in expectation. Each architecture we analyze in the experiments is trained $3$ times, with different parameter initialization and training samples: the SHD presented in the plots is the average of each of the $3$ models on $1500$ distinct test datasets of $1500$ points each, and the error bars are $95\%$ confidence intervals.